solve the folowing proportion for z
-15 /z = -30 /z -5

Answer by checkley77(12844)

-15/z=-30/(z-5)
-30Z=-15(Z-5)
-30Z=-15Z+75
-30Z+15Z=75
-15Z=75
Z=75/-15
Z=-5 ANS.
PROOF:
-15/-5=-30/(-5-5)
3=-30/-10
3=3